OpenStudy (anonymous):

An object is thrown upward from the top of a 192-foot building with an initial velocity of 16 feet per second. The height of the object after the t seconds is given by -16t^2+16+192. Factor this polynomial.

OpenStudy (anonymous):

the website wants the answer written a certain way........... it wants it to look like this -16(t- )(t+ )

OpenStudy (anonymous):

check that is it \[\Large -16t^2+16t+192 \] ???

OpenStudy (anonymous):

-16t^2+16+192. Factor this polynomial.

OpenStudy (anonymous):

i assume it should be what i wrote above because this is the only way you will get required format of answer i think you are missing t with 16 . take -16 common -16(t^2-t-12) can be written as -16(t^2-4t+3t-12) -16( t(t-4)+3(t-4)) -16((t-4)(t+3))

OpenStudy (anonymous):

it's says sorry that is not correct..... this is what it has underneath................. First factor out the greatest common factor ( if there is one other 1 or -1) then factor the remaining trinomial

OpenStudy (saifoo.khan):

-16(t^2-t-13) @sami-21

OpenStudy (anonymous):

saifoo that is not the way they want to answer........... they want the answer to -16(t- )(t+ )

OpenStudy (anonymous):

it should be if they want that way then you must know than there must be term with t which you are missing because (t+a)(t-b)=t^2-bt+at-ab

OpenStudy (anonymous):

How I wrote the question on here is the exact way that its it on the website

OpenStudy (anonymous):

lets try what you wrote can be written as -16(t^2-13) then use \[\Large a^2-b^2=(a+b)(a-b)\] \[\Large -16(t^2- \sqrt{13)}^2=-16(t+\sqrt{13})(t-\sqrt{13})\]

OpenStudy (anonymous):

no that is not the right answer either....... I get the same sorry that is not the correct answer

OpenStudy (anonymous):

Maby try multiplying to get rid of the "-" but i doubt thats it, and only do that if you have unlimited answers for your question

OpenStudy (anonymous):

post a screenie of the solution field(s)

OpenStudy (anonymous):

I have three chances to get the right answer then I have to go to another problem like it.. This is the 7 problem that I have worked.

OpenStudy (anonymous):

I can't do a screenie

hartnn (hartnn):

it says beneath that its a 'trinomial' means it must contain 3 different terms---> -16t^2+16t+192 -16t^2+16+192 is a binomial, not trinomial....

OpenStudy (anonymous):

it's a polyomial

OpenStudy (anonymous):

An object is thrown upward from the top of a 192-foot building with an initial velocity of 16 feet per second. The height of the object after the t seconds is given by -16t^2+16+192. Factor this polynomial.

hartnn (hartnn):

but is it explicitly mentioned beneath "factor the remaining trinomial" ? because what remians is binomial, not trinomial

OpenStudy (anonymous):

I have been on the problem for an hour and half now

OpenStudy (anonymous):

I have no ideal what you are talking about hartnn.......... I reposted the questions

OpenStudy (anonymous):

they want the answer to be in form of this -16(t- )(t+ )